OpenStudy (jellybot23):

What rectangular equation is equivalent to the polar equation r csc(theta)=8 Help pleaseeeeee

OpenStudy (jellybot23):

@Mehek14 @aaronq ??

OpenStudy (welshfella):



OpenStudy (welshfella):

csc

OpenStudy (welshfella):

csc theta = r / y = sqrt(x^2+ y^2) / y csc theta = 8/r = 8 / sqrt(x^2 + y^2) therefore sqrt(x^2+ y^2) / y = 8 / sqrt(x^2 + y^2)

OpenStudy (jellybot23):

how did you get √x^2+y^2 ?

OpenStudy (welshfella):

pythagoras theorem

OpenStudy (welshfella):

you can simplify the above by cross multiplying

OpenStudy (jellybot23):

Oh okay! Well, my choices are x=8, y=8, x=8y^2, and x^2+(y-4)^2=16 I am trying to cross multiply now and see if I can get it

OpenStudy (jellybot23):

I got x=8?

OpenStudy (jellybot23):

@welshfella

OpenStudy (welshfella):

No if you cross multiply my equation you get x^2 + y^2= 8x

OpenStudy (jellybot23):

Oh, I got x^2+y^2 = 8y

OpenStudy (welshfella):

if you simplify one of the choices you'll see that you get the above This choice will be the answer.

OpenStudy (welshfella):

OH sorry its 8y not 8x

OpenStudy (welshfella):

so - do you see which is the correct option?